OpenStudy (anonymous):

y is the midpoint of XZ . X Has Coordinates (2,4) , And Y has Coordinates (-1,1) . Find the coordinates of Z

OpenStudy (anonymous):

Using the Pythagorean Theorem, you should put the distance between Y and X as a triangle. So, starting by using the x-coordinate, Y is 3 units to the left of X. Therefore, one of the legs is 3 long. Now, the y-coordinate. Y is 3 below X, so 3 is also the other leg. So, the Pythagorean Theorem is 3^2 + 3^2 = c^2 (c is the hypotenuse, or missing leg). You should receive the result of c^2 = 18, or c = 3√2. Since c is the distance between X and Y, c is also the distance between Y and Z since Y is the midpoint of X and Z. The Answer: 3√2

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank You Now I Understand. It Was Confusing, But Now I Get It . Thanks !